Six drivers were traveling on the highway at the rates listed below.How many miles per hour was each driver traveling? How do you right each answer from slowest to the fastset?
1. Dan: 110 miles in 2 hours 4.Nan :300 miles in 5 hours
2.Stan: 204 miles in 3 hours 5.Jan :14 miles in 15 minutes
3.Ann: 25 miles in 1/2 hour 6.Fran:22 miles in 20 minutes

Answer by Mathtut(3670)

mph is the number of miles driven divided by the time in hours
:
1. Dan: 110 miles in 2 hours 4.Nan :300 miles in 5 hours
2.Stan: 204 miles in 3 hours 5.Jan :14 miles in 15 minutes
3.Ann: 25 miles in 1/2 hour 6.Fran:22 miles in 20 minutes
:
for instance Dan would be: 110/2=55 mph
:
Nan:300/5=60 mph
stan: 204/3=68 mph
:
I will let you do the rest.....remember if its in minutes you have to convert to hours so 20 minutes is 1/3 hour..etc
:
I will do one fractional hour problem
:Ann:25/(1/2)=50 mph.....inverted 1/2 and multiplied
:
have fun you have 2 left